# ResGS: Residual Densification of 3D Gaussian for Efficient Detail Recovery

### v2

@@ 4.2 Comparisons
To assess the effectiveness of our method, we compared it with several 3D-GS variants: 3D-GS [24], Scaffold-GS [29], Octree-GS [35], FreGS [48], AbsGS [43], Pixel-GS [50] and Mini-Splatting-D [13]. Note that Scaffold-GS and Octree-GS use a latent representation rather than the original 3D-GS representation. Some radiance-field-based methods are also selected for comparison, including Plenoxels [45], Instant-NGP [32] and Mip-NeRF360 [1]. We found that the resolution of Mip-NeRF360 dataset images varies across different works. Following the resolution setup in the 3D-GS work, we retrain some models on the Mip-NeRF360 dataset. Additionally, for works that did not report their storage, we obtained and reported it based on their released code.
@@
In Table 1, we present quantitative results comparing the rendering quality of our method with other approaches. From the table, we can see that our method achieves SOTA performance across all metrics on the Mip-NeRF360 [1] dataset. On the Tanks&Temples [25] dataset, we also achieve the best SSIM and LPIPS metrics, which closely reflect human perception. In the Deep Blending [18] dataset, while our method does not achieve the highest score, it shows a notable improvement over the baseline. Additionally, our model variant without AbsGS [43] maintains similar performance to the full model, with only a slight reduction in LPIPS. Notably, our smaller model version achieves strong performance with reduced memory usage, showcasing the efficiency and scalability of our approach. In Fig. 5, we present qualitative comparisons between our method and several 3D-GS variants [24, 50, 43] across various scenes, including both indoor and outdoor settings. It is evident that our model captures finer details, as shown in the second, third, and fourth rows, constructs missing geometry more effectively (last row), and achieves more accurate geometry reconstruction (first row).
@@
Here, we provide the computational costs of our method in comparison with other approaches. The results are shown in Table. 2. From the table, we can see that our method involves a slight increase in training time while delivering much higher rendering quality than methods with lower rendering quality like 3D-GS [24] and AbsGS [43]. Our method is also significantly more efficient than other methods like Mini-Splatting-D [13], with comparable rendering quality. Notably, our small model provides the fastest FPS and shortest training time in all compared methods while maintaining high fidelity.
@@ 4.3.1 Residual Split
We analyze the compatibility of our novel densification operation by replacing split and clone in multiple 3D-GS variants [24, 43, 50, 13] with residual split. The hyperparameters were kept the same as in the original pipelines. The results are shown in Table 3. The table shows that our method leads to a notable increase in PSNR metrics across all pipelines, while for other metrics, some exhibit improvements, and others remain approximately the same. Additionally, memory consumption is reduced after incorporating residual split compared to the original pipeline. This underscores our method’s ability to improve rendering quality and reduce redundancy, while also demonstrating its compatibility and potential for broader application across 3D-GS-oriented works as an effective alternative to split and clone.
@@
Since residual split involves some hyperparameters: opacity reduction weight $\beta$ and scale reduction factor $\lambda_{s}$, we conducted a corresponding sensitivity analysis. We tested our method upon the Tanks&Temples dataset while applying different $\beta\in[0.05,1.0)$ and $\lambda_{s}\in[1.05,2.0]$. All the models are trained with similar storage. The LPIPS scores are shown in Fig. 6 and 7. As shown in the figures, LPIPS scores remain relatively stable when $\beta<0.4$ and $\lambda_{s}>1.55$. These results indicate that our method is robust to hyperparameter variations, with only a 0.006 difference in LPIPS between the best and worst cases, and a relatively broad range where rendering quality remains optimal.
@@ 4.3.2 Ablation Studies
Table 4: Ablation Study on our pipeline. Here, IP denotes image pyramid supervision, RS represents using residual split for densification, and VT indicates applying varying gradient threshold. Each component contributes to our method’s effectiveness.
Dataset Mip-NeRF360
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Base 27.41 0.817 0.189
Base + RS 27.66 0.825 0.183
Base + IP 27.54 0.823 0.182
Base + RS + IP 27.88 0.831 0.178
Base + RS + IP + VT (full) 28.00 0.833 0.174
To assess the effectiveness of individual modules in our proposed pipeline—residual split, image pyramid as supervision, and varying gradient threshold—ablation studies on the Mip-NeRF360 [1] dataset were conducted. For the varying gradient threshold, we only conducted studies on its removal, as it is highly interdependent with the other two modules. The results are shown in Table 4. It is evident that both residual split and the image pyramid supervision individually enhance the quality of our baseline. However, we observe that adding image pyramid supervision alone does not lead to significant improvements. This is because the split-and-clone operations do not perform well within a coarse-to-fine schedule. In the early stages, when the overall structure is still being constructed, split is primarily chosen, which exacerbates the issue rather than improving it. On the other hand, in later iterations, when fine details need to be recovered, clone is selected and does not address the issue properly. Conversely, residual split works perfectly due to its adaptiveness. This is evident from the notable improvement of Base + RS + IP over both Base + IP and Base + RS, which also shows that a coarse-to-fine schedule supplements residual split as expected. Additionally, applying the varying gradient threshold further boosts performance. In conclusion, each module contributes to improved results.
@@ 4.4 Discussions and Limitations
Our results show that our method significantly improves rendering quality on the Mip-NeRF360 [1] and Tanks&Temples [25] datasets. However, for Deep Blending [18], our method did not yield a significant boost in performance. This variation is primarily due to the challenges specific to each dataset: Mip-NeRF360 and Tanks&Temples require capturing fine textures and recovering missing geometry, whereas Deep Blending focuses on handling occlusions and reconstructing accurate geometry in weakly textured areas. Our method mainly contributes to retrieving fine details and sufficient geometry and does not address the issue of occlusion and weak-textured areas, suggesting it to be a topic for future optimization.
@@ 5 Conclusion
In this paper, we analyzed the limitations of the current 3D-GS densification operation. To address these issues, we propose residual split, a residual empowered densification operation that adaptively handles both over-reconstruction and under-reconstruction. Additionally, we introduce ResGS, a coarse-to-fine 3D-GS pipeline to supplement residual split. Specifically, we use an image pyramid for supervision to gradually recover the overall structure to fine details during Gaussian optimization. We also design a Gaussian selection scheme to further densify coarse Gaussians progressively. Experiments show our pipeline achieves SOTA rendering quality while residual split is highly compatible with other 3D-GS variants.
@@ 6.2 Further Analysis of Residual Split on 3D-GS
Table 9: Results of our method without extended periodic opacity reduction. Here, No EPOR refers to our method trained without applying the extended periodeic opacity reduction. No EPOR + CT refers to the model without the extended periodic opacity reduction but with an adjusted gradient threshold to ensure that the final storage size is approximately the same as our full model with extended periodic opacity reduction.
@@ 6.4 Image Resolution of Mip-NeRF360 Dataset
In our paper, we mentioned that the image resolution on the Mip-NeRF360 [1] dataset varies across different works. 3D-GS [24] uses the provided “images_4”, the official downscaled 4 times images from the dataset, for outdoor scenes and “images_2” for indoor scenes. Most works [43, 13, 48] follow the 3D-GS setting, but Scaffold-GS [29], Octree-GS [35], Pixel-GS [50] uses different settings. Specifically, Scaffold-GS and Octree-GS downscale original images to 1.6K resolution, while Pixel-GS downscales original images by a factor of 4 for outdoor scenes and 2 for indoor scenes during the training process, instead of using the provided downscaled images. This discrepancy results in differences in experimental settings. To further demonstrate the performance of our method, we provide its results under the two different settings, as shown in Table 10 and 11. The tables show that our method prevails even under these alternate settings.
@@ 6.5 Per Scene Results
We present the per-scene results of the used metrics in Table 12-17. For works that did not report per-scene results, we obtained and reported them using their released code. The works selected for comparison are mainly the same as in the paper: 3D-GS [24], Scaffold-GS [29], Octree-GS [35], AbsGS [43], Pixel-GS [50], Mini-Splatting-D [13], Plenoxels [45], Instant-NGP [32] and Mip-NeRF360 [1]. Note that we did not show the results of FreGS [48] since they did not report their per-scene results or release their code. Works retrained on the Mip-NeRF360 [1] are marked with a .
@@
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 23.22 18.93 23.14 22.98
Instant-NGP 23.38 20.46 28.26 21.67
Mip-NeRF360 24.91 19.52 29.14 29.66
Scaffold-GS 25.77 22.15 29.80 30.62
Octree-GS 26.27 22.77 29.87 30.95
3D-GS 25.19 21.10 28.77 30.04
AbsGS 25.74 21.72 29.20 30.14
Pixel-GS 25.49 22.13 28.02 29.79
Mini-Splatting-D 25.43 21.04 29.32 30.43
Ours-Small (AbsGS) 26.07 22.40 29.64 30.38
Ours (3D-GS) 26.08 22.65 29.54 30.32
Ours (AbsGS) 26.14 22.61 29.51 30.32
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.774 0.663 0.787 0.802
Instant-NGP 0.800 0.689 0.854 0.779
Mip-NeRF360 0.857 0.660 0.901 0.900
Scaffold-GS 0.883 0.822 0.907 0.904
Octree-GS 0.896 0.835 0.912 0.914
3D-GS 0.879 0.802 0.899 0.906
AbsGS 0.888 0.818 0.898 0.907
Pixel-GS 0.883 0.823 0.885 0.899
Mini-Splatting-D 0.890 0.817 0.905 0.908
Ours-Small (AbsGS) 0.889 0.834 0.905 0.907
Ours (3D-GS) 0.890 0.840 0.901 0.905
Ours (AbsGS) 0.893 0.841 0.900 0.905
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.335 0.422 0.521 0.499
Instant-NGP 0.249 0.360 0.352 0.428
Mip-NeRF360 0.159 0.354 0.237 0.252
Scaffold-GS 0.147 0.206 0.250 0.258
Octree-GS 0.118 0.198 0.231 0.244
3D-GS 0.148 0.218 0.244 0.241
AbsGS 0.132 0.193 0.241 0.233
Pixel-GS 0.122 0.180 0.258 0.243
Mini-Splatting-D 0.100 0.181 0.218 0.204
Ours-Small (AbsGS) 0.124 0.177 0.233 0.235
Ours (3D-GS) 0.118 0.164 0.232 0.231
Ours (AbsGS) 0.106 0.159 0.231 0.223

### v3

@@ 4.2 Comparisons
To assess the effectiveness of our method, we compare it with several 3D-GS variants: 3D-GS [23], Scaffold-GS [27], Octree-GS [33], FreGS [48], AbsGS [43], Pixel-GS [50] and Mini-Splatting-D [12]. Note that Scaffold-GS and Octree-GS use a latent representation rather than the original 3D-GS representation. Some radiance-field-based methods are also selected for comparison, including Plenoxels [45], Instant-NGP [30] and Mip-NeRF360 [1]. We find that the resolution of Mip-NeRF360 dataset images varies across different works. Following the resolution setup in the 3D-GS work, we retrain some models on the Mip-NeRF360 dataset. Additionally, for works that did not report their storage usage, we report the results based on their released code.
@@
In Table 1, we present quantitative results comparing the rendering quality of our method with other approaches. From the table, we can see that our method achieves SOTA performance across all metrics on the Mip-NeRF360 [1] dataset. On the Tanks&Temples [24] dataset, our method also achieves the best SSIM and LPIPS metrics, which closely reflect human perception. In the Deep Blending [17] dataset, though our method does not achieve the highest score, it shows a notable improvement over the baseline. Additionally, our model variant without AbsGS [43] maintains similar performance to the full model, with only a slight reduction in LPIPS. Notably, our smaller model version achieves strong performance with reduced memory usage, showcasing the efficiency and scalability of our approach. In Fig. 5, we present qualitative comparisons between our method and several 3D-GS variants [23, 50, 43] across various scenes, including both indoor and outdoor settings. It is evident that our model captures finer details, as shown in the second, third, and fourth rows, constructs missing geometry more effectively (last row), and achieves more accurate geometry reconstruction (first row).
@@
We provide the computational costs of our method in comparison with other approaches in Table 2. It shows that our method involves a slight increase in training time while delivering much higher rendering quality than methods with lower rendering quality like 3D-GS [23] and AbsGS [43]. Our method is also significantly more efficient than other methods like Mini-Splatting-D [12], with comparable rendering quality. Notably, our small model provides the fastest FPS and shortest training time in all compared methods while maintaining high fidelity.
@@ 4.3.1 Residual Split
We analyze the compatibility of our novel densification operation by replacing split and clone in multiple 3D-GS variants [23, 43, 50, 12] with residual split. The hyperparameters are set the same as in the original pipelines. As Table 3 shows, our method leads to a notable increase in PSNR metrics across all pipelines, while for other metrics, some exhibit improvements, and others remain approximately the same. Additionally, memory consumption is reduced after incorporating residual split compared to the original pipeline. This underscores our method’s ability to improve rendering quality and reduce redundancy, while also demonstrating its compatibility and potential for broader application across 3D-GS-oriented works as an effective alternative to split and clone.
@@
Our residual split involves some hyperparameters: opacity reduction weight $\beta$ and scale reduction factor $\lambda_{s}$. We conduct a corresponding sensitivity analysis on the Tanks&Temples [24] dataset while applying different $\beta\in[0.05,1.0)$ and $\lambda_{s}\in[1.05,2.0]$. All models are trained with similar storage. The LPIPS scores with different parameters are shown in Fig. 6 and 7. As they show, the LPIPS score remains relatively stable when $\beta<0.4$ and $\lambda_{s}>1.55$, indicating that our method is robust to hyperparameter variations, with only a 0.006 difference in LPIPS between the best and worst cases and a relatively broad range where rendering quality remains optimal.
@@ 4.3.2 Ablation Studies
Table 4: Ablation Study on our pipeline. IP denotes image pyramid supervision, RS represents using residual split for densification, and VT indicates applying varying gradient threshold. Each component contributes to our method’s effectiveness.
Dataset Mip-NeRF360 Deep Blending
@@
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Base 27.41 0.817 0.189 29.60 0.899 0.240
Base + RS 27.66 0.825 0.183 29.68 0.900 0.233
Base + IP 27.54 0.823 0.182 29.05 0.896 0.245
Base + RS + IP 27.88 0.831 0.178 29.82 0.902 0.230
Base + RS + IP + VT (full) 28.00 0.833 0.174 29.91 0.902 0.227
To assess the effectiveness of our pipeline’s key modules, including residual split, image pyramid as supervision, and the varying gradient threshold, we conduct ablation studies on the Mip-NeRF360 [1] and Deep Blending [17] datasets. For the varying gradient threshold, we only conduct studies on its removal, as it is highly interdependent with the other two modules. As the results in Table 4 shows, both residual split and image pyramid supervision individually boost the baseline’s performance on the Mip-NeRF360 dataset.
However, on Deep Blending, we observe a performance drop in Base+IP. This is because the dataset contains many weakly textured areas. As noted in Sec. 1, split and clone tend to fill such regions with an excessive number of small Gaussians. This makes them prone to overfitting the training views and lead to reduced performance on test views. The issue is amplified in the early iterations of a coarse-to-fine schedule, where only coarse scene information is available, making weakly textured areas even harder to resolve. As a result, these small Gaussians struggle to construct accurate geometry, leading to the performance drop observed in Base+IP.
Conversely, when using RS, weakly textured areas in early iterations are represented by only a few Gaussians, reducing overfitting. In later iterations, as finer information becomes available, RS progressively introduces more small Gaussians. As a result, RS benefits from progressive training, as evidenced by the improved performance of Base+RS+IP over Base+RS. This further proves RS is more effective than split and clone, especially in a coarse-to-fine schedule. Additionally, applying the varying gradient threshold (our full model) further boosts performance on both datasets, validating its effectiveness.
@@ 4.4 Discussions and Limitations
The above results demonstrate that our method significantly enhances rendering quality on the Mip-NeRF360 [1] and Tanks&Temples [24] datasets, achieving state-of-the-art performance. However, for the Deep Blending dataset [17], our method fails to achieve the best performance on every metric. While our approach implicitly constrains the number of Gaussians in weakly textured regions, which helps mitigate overfitting and improves quality, it does not include explicit regularization. This suggests an opportunity for further optimization in future work.
@@ 5 Conclusion
In this paper, we analyze the limitations of the current 3D-GS densification operation. To address these issues, we propose residual split, a residual empowered densification operation that adaptively handles both over-reconstruction and under-reconstruction. Additionally, we introduce ResGS, a coarse-to-fine 3D-GS pipeline to supplement residual split. Specifically, we use an image pyramid for supervision to gradually introduce fine details during Gaussian optimization. We also design a Gaussian selection scheme to further densify coarse Gaussians progressively. Our pipeline achieves SOTA rendering quality on several datasets while our residual split is highly compatible with other 3D-GS variants.
Acknowledgements
This work was supported by CAS Project for Young Scientists in Basic Research (YSBR-116).
@@ 6 Additional Experiment Results
6.1 Evaluation on large-scale urban datasets
To evaluate the performance of our approach on large-scale urban datasets, we performed experiments using the Waymo [36] dataset and compared our results against GaussianPro [11]. The results are shown in the table below. It can be seen that our method achieves better results across all metrics.
Table 5: Results on the Waymo dataset compared with GaussianPro.
Dataset Waymo
Method Metrics
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Mem.
GaussianPro 34.68 0.949 0.191 297MB
Ours 35.29 0.951 0.191 284MB
@@ 6.3 Further Analysis of Residual Split on 3D-GS
Table 10: Results of our method without extended periodic opacity reduction. Here, No EPOR refers to our method trained without applying the extended periodic opacity reduction. No EPOR + CT refers to the model without the extended periodic opacity reduction but with an adjusted gradient threshold to ensure that the final storage size is approximately the same as our full model with extended periodic opacity reduction.
@@
6.4 Results using random initialization
To assess the robustness of our method to random initialization, we conducted experiments using randomly initialized point clouds and compared our approach with other methods trained on the same random point clouds. The results are shown in Table 11. For the Mip-NeRF360 [1] and Tanks&Temples [24] datasets, our model shows a noticeable drop in quality compared to when initialized with SfM [35]. However, compared to other methods, it still achieves the best performance across all three metrics on Mip-NeRF360, and the highest SSIM and LPIPS on Tanks&Temples. As for Deep Blending [17], our method and others, excluding Scaffold-GS and Octree-GS, do not exhibit a significant performance drop. While we do not achieve the highest scores, our method remains competitive, ranking second across all metrics.
Table 11: Results using random initialized point clouds, compared with other methods.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method Metrics
Scaffold-GS 25.90 0.743 0.301 157MB 22.00 0.768 0.282 51MB 29.48 0.891 0.285 56MB
Octree-GS 25.70 0.739 0.292 190MB 22.15 0.794 0.226 81MB 28.29 0.873 0.300 92MB
3D-GS 25.90 0.769 0.266 632MB 21.19 0.777 0.241 339MB 29.42 0.897 0.253 605MB
AbsGS 25.50 0.773 0.226 741MB 20.89 0.786 0.211 304MB 29.52 0.897 0.241 478MB
Mini-Splatting-D 25.18 0.792 0.207 1.10GB 19.77 0.760 0.220 1.02GB 29.76 0.901 0.212 1.09GB
Pixel-GS 26.23 0.792 0.221 1.14GB 21.10 0.788 0.207 939MB 28.76 0.889 0.256 1.04GB
Ours-Random 26.38 0.799 0.200 750MB 22.02 0.815 0.178 282MB 29.57 0.899 0.234 464MB
Ours-SfM 28.00 0.833 0.174 698MB 24.38 0.867 0.132 351MB 29.91 0.902 0.227 586MB
@@ 6.6 Image Resolution of Mip-NeRF360 Dataset
In our paper, we mentioned that the image resolution on the Mip-NeRF360 [1] dataset varies across different works. 3D-GS [23] uses the provided “images_4”, the official downscaled 4 times images from the dataset, for outdoor scenes and “images_2” for indoor scenes. Most works [43, 12, 48] follow the 3D-GS setting, but Scaffold-GS [27], Octree-GS [33], Pixel-GS [50] uses different settings. Specifically, Scaffold-GS and Octree-GS downscale original images to 1.6K resolution, while Pixel-GS downscales original images by a factor of 4 for outdoor scenes and 2 for indoor scenes during the training process, instead of using the provided downscaled images. This discrepancy results in differences in experimental settings. To further demonstrate the performance of our method, we provide its results under the two different settings, as shown in Table 12 and 13. The tables show that our method prevails even under these alternate settings.
@@ 6.7 Per Scene Results
We present the per-scene results of the used metrics in Table 14-19. For works that did not report per-scene results, we obtained and reported them using their released code. The works selected for comparison are mainly the same as in the paper: 3D-GS [23], Scaffold-GS [27], Octree-GS [33], AbsGS [43], Pixel-GS [50], Mini-Splatting-D [12], Plenoxels [45], Instant-NGP [30] and Mip-NeRF360 [1]. Note that we did not show the results of FreGS [48] since they did not report their per-scene results or release their code. Works retrained on the Mip-NeRF360 [1] are marked with a .
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 23.22 18.93 23.14 22.98
Instant-NGP 23.38 20.46 28.26 21.67
Mip-NeRF360 24.91 19.52 29.14 29.66
Scaffold-GS 25.77 22.15 29.80 30.62
Octree-GS 26.27 22.77 29.87 30.95
3D-GS 25.19 21.10 28.77 30.04
AbsGS 25.74 21.72 29.20 30.14
Pixel-GS 25.49 22.13 28.02 29.79
Mini-Splatting-D 25.43 21.04 29.32 30.43
Ours-Small (AbsGS) 26.07 22.40 29.64 30.38
Ours (3D-GS) 26.08 22.65 29.54 30.32
Ours (AbsGS) 26.14 22.61 29.51 30.32
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.774 0.663 0.787 0.802
Instant-NGP 0.800 0.689 0.854 0.779
Mip-NeRF360 0.857 0.660 0.901 0.900
Scaffold-GS 0.883 0.822 0.907 0.904
Octree-GS 0.896 0.835 0.912 0.914
3D-GS 0.879 0.802 0.899 0.906
AbsGS 0.888 0.818 0.898 0.907
Pixel-GS 0.883 0.823 0.885 0.899
Mini-Splatting-D 0.890 0.817 0.905 0.908
Ours-Small (AbsGS) 0.889 0.834 0.905 0.907
Ours (3D-GS) 0.890 0.840 0.901 0.905
Ours (AbsGS) 0.893 0.841 0.900 0.905
Method Scenes
Truck Train Dr Johnson Playroom
Plenoxels 0.335 0.422 0.521 0.499
Instant-NGP 0.249 0.360 0.352 0.428
Mip-NeRF360 0.159 0.354 0.237 0.252
Scaffold-GS 0.147 0.206 0.250 0.258
Octree-GS 0.118 0.198 0.231 0.244
3D-GS 0.148 0.218 0.244 0.241
AbsGS 0.132 0.193 0.241 0.233
Pixel-GS 0.122 0.180 0.258 0.243
Mini-Splatting-D 0.100 0.181 0.218 0.204
Ours-Small (AbsGS) 0.124 0.177 0.233 0.235
Ours (3D-GS) 0.118 0.164 0.232 0.231
Ours (AbsGS) 0.106 0.159 0.231 0.223